### v1

@@ 5 Conclusion
We explore the role of AGN feedback on the stellar kinematics of galaxies based on EAGLE simulations with various AGN feedback models. Our analysis is based on five EAGLE runs—EAGLE-50, EAGLE-eAGN-dT, EAGLE-eAGN-Visc, EAGLE-wAGN, and EAGLE-NoAGN—each covering a cubic volume of (50 Mpc)3. We select pressure-dominated galaxies based on stellar kinematics ($v/\sigma<0.5$) for log ($M_{*}/\rm{M}_{\odot})>9$. Furthermore, we construct the velocity dispersion complete sample in five EAGLE simulations. We then measure stellar velocity dispersions using the same approach we applied for fiber spectroscopy.
@@
The stellar mass–velocity dispersion relations derived from the EAGLE simulations show a significant offset compared to observations: simulated galaxies exhibit systematically lower velocity dispersions at fixed stellar mass. This trend is consistent with results from other cosmological simulations, including IllustrisTNG-100 and 300 (Sohn2024a). As a result, the velocity dispersion functions (VDFs) derived from all five EAGLE runs are significantly lower than the observed VDF.
@@
Despite the systematic offset, comparing the stellar velocity dispersions across EAGLE models with different AGN feedback implementations reveals the impact of AGN feedback on stellar kinematics. The VDF from EAGLE-NoAGN shows a substantial excess of high velocity dispersion galaxies ($\log\sigma_{*}>2.3$) and a flatter slope at $\log\sigma_{*}>2.0$ relative to those from other EAGLE simulations. This discrepancy strongly suggests that the AGN feedback is certainly responsible for the stellar velocity dispersion (and the underlying mass) distribution of galaxies.
@@
In order to investigate the physical mechanisms of AGN feedback that affect the stellar kinematics, we investigate the $M_{*}-\sigma_{*}$ scaling relations measured within a fixed 10 kpc aperture in EAGLE simulations. The $M_{*}-\sigma_{*}$ relations in all EAGLE simulations have essentially identical shapes that velocity dispersions are simply proportional to the stellar mass enclosed in the aperture we use for measuring velocity dispersion. In other words, the discrepancy in the VDFs results from the underlying mass distribution rather than any peculiarity in the scaling relation between mass and stellar kinematics.
@@
We further show that stellar mass concentration varies significantly with AGN feedback. Galaxies in EAGLE-NoAGN display much stronger central mass concentrations (within 10 kpc) than their counterparts in the other runs. This is primarily driven by sustained central star formation in the absence of AGN feedback, as opposed to the suppressed central star formation seen in simulations with AGN. These differences in stellar mass concentration naturally propagate to differences in stellar velocity dispersions.
@@
The VDF offers a promising probe of the dark matter (DM) subhalo mass function (Sohn2017b; Sohn2024b). Prior studies (Zahid2018; Sohn2024a) have highlighted a tight correlation between stellar velocity dispersion and the velocity dispersion of the host DM halo. As a next step, comparing the DM mass functions and DM VDFs from simulations will help place further constraints on the stellar VDF as a tracer of the underlying dark matter distribution.
@@ Acknowledgements.
We acknowledge the Virgo Consortium for making their simulation data available. The EAGLE simulations were performed using the DiRAC-2 facility at Durham, managed by the ICC, and the PRACE facility Curie based in France at TGCC, CEA, Bruyères-le-Châtel. J.S. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2023-00210597). This work was also supported by the Global-LAMP Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2023-00301976).

### v2

@@ 5 Conclusion
We explored the role of AGN feedback in the stellar kinematics of galaxies based on EAGLE simulations with various AGN feedback models. Our analysis is based on five EAGLE runs (EAGLE-50, EAGLE-eAGN-dT, EAGLE-eAGN-Visc, EAGLE-wAGN, and EAGLE-NoAGN) each covering a cubic volume of (50 Mpc)3. We selected pressure-dominated galaxies based on stellar kinematics ($v/\sigma<0.5$) for log ($M_{*}/\rm{M}_{\odot})>9$. Furthermore, we constructed the velocity dispersion complete sample in five EAGLE simulations. We then measured stellar velocity dispersions using the same approach we applied for fiber spectroscopy.
@@
The stellar mass-velocity dispersion relations derived from the EAGLE simulations show a significant offset compared to observations: simulated galaxies exhibit systematically lower velocity dispersions at fixed stellar mass. This trend is consistent with results from other cosmological simulations, including IllustrisTNG-100 and 300 (Sohn et al., 2024b). As a result, the velocity dispersion functions (VDFs) derived from all five EAGLE runs are significantly lower than the observed VDF.
@@
Despite the systematic offset, comparing the stellar velocity dispersions across EAGLE models with different AGN feedback implementations reveals the impact of AGN feedback on stellar kinematics. The VDF from EAGLE-NoAGN shows a substantial excess of high-velocity dispersion galaxies ($\log\sigma_{*}>2.3$) and a flatter slope at $\log\sigma_{*}>2.0$ relative to those from other EAGLE simulations. This discrepancy strongly suggests that the AGN feedback is certainly responsible for the stellar velocity dispersion (and the underlying mass) distribution of galaxies.
@@
In order to investigate the physical mechanisms of AGN feedback that affect the stellar kinematics, we investigated the $M_{*}-\sigma_{*}$ scaling relations measured within a fixed 10 kpc aperture in EAGLE simulations. The $M_{*}-\sigma_{*}$ relations in all EAGLE simulations have essentially identical shapes; velocity dispersions are simply proportional to the stellar mass enclosed in the aperture we used for measuring velocity dispersion. In other words, the discrepancy in the VDFs results from the underlying mass distribution rather than any peculiarity in the scaling relation between mass and stellar kinematics.
@@
We further show that stellar mass concentration varies significantly with AGN feedback. Galaxies in EAGLE-NoAGN display much stronger central mass concentrations (within 10 kpc) than their counterparts in the other runs. This is primarily driven by sustained central star formation in the absence of AGN feedback, as opposed to the suppressed central star formation seen in simulations with AGNs. These differences in stellar mass concentration naturally propagate to differences in stellar velocity dispersions.
@@
The VDF offers a promising probe of the DM subhalo mass function (Sohn et al., 2017b, 2024a). Prior studies (Zahid et al., 2018; Sohn et al., 2024b) have highlighted a tight correlation between stellar velocity dispersion and the velocity dispersion of the host DM halo. As a next step, comparing the DM mass functions and DM VDFs from simulations will help place further constraints on the stellar VDF as a tracer of the underlying DM distribution.
@@ Acknowledgements.
We thank the referee for carefully reading our manuscript and for the helpful comments that significantly improved it. We acknowledge the Virgo Consortium for making their simulation data available. The EAGLE simulations were performed using the DiRAC-2 facility at Durham, managed by the ICC, and the PRACE facility Curie based in France at TGCC, CEA, Bruyères-le-Châtel. J.S. is supported by the National Research Foundation of Korea (NRF) grant funded by the Korean government (MSIT) (RS-2023-00210597). This work was also supported by the Global-LAMP Program of the National Research Foundation of Korea (NRF) grant funded by the Ministry of Education (No. RS-2023-00301976).